# High-Capacity Silicon Microdisk Transmitter

### v1

@@ Abstract
Optical interconnects are the most promising solution to address the data-movement bottleneck in data centers. Silicon microdisks, benefiting from their compact footprint, low energy consumption, and wavelength division multiplexing (WDM) capability, have emerged as an attractive and scalable platform for optical modulation. However, microdisk resonators inherently exhibit low fabrication error tolerance, limiting their practical deployment. Here, utilizing a CMOS photonics platform, we demonstrate 1.2 Tb/s of off-die bandwidth through a 64 microdisk modulator system. In addition, we develop an automated, close-looped, non-reversible, low-loss, and picometer-precision permanent wavelength tuning technique using laser trimming. The trimming technique reduces 33 % of the energy consumption needed to thermally tune the microdisk resonant wavelength. Using this technique, we achieve a fully passive, 5-channel dense wavelength division multiplexing (DWDM, 50 GHz spacing) transmitter. The integration of the high speed (1.2 Tb/s), low energy consumption (29 fJ/bit) and the permanent wavelength trimming lays a robust foundation for next-generation optical interconnect systems, poised to facilitate scaling of future AI and computing hardware.
@@ Introduction
The exponential growth of artificial intelligence (AI) and deep learning applications drives computational workloads from individual devices to server racks in data centers, placing unprecedented requirements on the interconnect and communication infrastructures, which need to be high-speed and low-energy-consumption [1, 2, 3, 4, 5, 6, 7, 8]. Traditional intra- and inter- chip interconnects, which rely on electrical metal-wire links, are struggling to satisfy the increasing bandwidth and energy consumption requirements imposed by current data-hungry AI applications [3, 9, 2]. As AI workloads scale, multi-node, multi-processor systems demand significantly much higher-speed and more energy-efficient XPU-to-XPU data communications to effectively handle large-scale computational-intensive tasks [9, 10, 11, 12]. In contrast, optical interconnect technologies that leverage silicon photonics (SiPh) offer bandwidth and energy consumption advantages over conventional electrical metal-wire links [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. SiPh components, which are based on the total internal reflection of silicon-on-insulator (SOI) waveguides, have several distinct advantages, including compatibility with mature complementary metal–oxide–semiconductor (CMOS) manufacturing process, large bandwidth (10’s THz), compact footprint ($\leavevmode\nobreak\ \upmu\text{m}$ scale), low loss (dB/m), and intrinsic high parallelism [27, 28, 29, 30, 31, 32, 33]. These combined advantages make SiPh interconnects an essential infrastructure components to address the large-bandwidth and low-energy-consumption requirements for contemporary data center.
@@
For short-reach interconnects and communications, the low-complexity and low-cost intensity-modulated direct-detection (IM-DD) system remains dominant within the data centers [34, 16, 17], where integrated electro-optic intensity modulators (EOIMs) convert the XPU electrical signals to analog optical intensities. The high-index contrast of the SiPh platform enables the realization of compact resonant EOIMs which minimize the device footprint and capacitance: thus enabling high-speed operation, low energy consumption, and large-scale dense integration. In addition, by leveraging the intrinsic wavelength selectivity and compact footprint nature of these resonant structure, WDM and spatial division multiplexing (SDM) can be implemented to provide a scalable parallel and distinguishable high capacity modulation across a broadband wavelength range. Consequently, SiPh resonator EOIMs enables large-scale, high bandwidth, and low energy consumption modulation in a compact footprint [29, 27, 19, 17, 35]. The CMOS compatibility of SiPh also allows high-yield mass production, which is desired for low-cost optical interconnects and communications in data centers.
@@
Figure 1: Artistic vision and microscope image of the high parallelism, high capacity micro-disk transmitter system. (a) Schematic diagram of the 64-channel high parallelism, high capacity microdisk transmitter. The transmitter supports high parallelism SDM and allows for high per-fiber bandwidth WDM transmission. (b) Optical microscope image of the transmitter chip, from top to bottom, the chip includes the input edger coupler, splitting tree, phase shifters, microdisk modulators, 2D emitter array, and the 1D emitter array. (c) High resolution microscope images of the main components of the chip. (d) 2D emission image of the chip shows an 8 by 8 emitter array.
@@
Figure 2: Direct current characterization of the Microdisk Transmitter chip. (a) Experimental setup of the microdisk transmitter direct current characterization, the TLS and OSA are synchronized to provide high precision characterization, an IR camera is employed to provide parallel characterization. TLS, tunable laser; OSA, optical spectrum analyzer; PC, polarization controller; VGFA, V-groove fiber array; EDFA, Er-doped fiber amplifier. (b) Measured and fitted grating coupler transmission spectrum of the transmitter. (c) Measured broadband transmission spectrum of a single-channel transmitter at 0-V D.C voltage. (d) Measured transmission spectrum versus reversed junction voltage from 0.5 V to -1.5 V, under a 2-V voltage swing at 1536.38 nm, the modulator insertion loss (IL) is 2.6 dB, the extinction ratio (ER) is 9.8 dB, and the optical modulation amplitude (OMA) is 0.49, these parameter values are nearly constant among these 64-channel microdisk modulators. (e) Measured transmission spectrum versus heater voltage ranging from 0 V to 5 V, showing a wavelength red shift tunning efficiency of 420 pm/mW. (f) Extracted absorption coefficient and refractive index change of the microdisk transmitter versus junction voltages from 0.5 V to -1.5 V.
@@
While microdisk EOIMs offer compact footprints, they exhibit extreme sensitivity to fabrication errors and uncertainties [35, 36, 37, 38]. Even sub-nanometer variations in waveguide dimensions can significantly shift the spectral response of these devices. Both active tuning and post-fabrication trimming methods are employed to correct the resonance uncertainties [39, 38, 40, 41]. Active tuning, commonly achieved through the thermo-optic effect, though effective, it needs continuously large power consumption. In contrast, post-fabrication trimming permanently modify the effective index of the optical mode and doesn’t need continuous power consumption. Established techniques include selective post-fabrication etching, deposition of phase change materials (PCMs), anodic oxidation, laser and thermal annealing [38, 42, 43, 44]. However, in standard foundry-fabricated silicon modulators the waveguides are buried beneath thick dielectric top claddings—and thus less compatible with PCM integration—criss-crossed by embedded metal interconnects, and incorporate active p–n junctions that cannot tolerate the elevated post-process temperatures required by some trimming schemes. Instead, adopt a method to locally densify the oxide cladding of each modulator via laser annealing as a practical post-fabrication trimming technique [45]. We observe that this process increases the mode effective refractive index of the microdisk EOIM, yielding a red shift in the resonant wavelength.
@@
In this work, we demonstrate and perform non-volatile trimming of a highly parallel, high capacity EOIM transmitter featuring 64 channels compact microdisks integrated onto a single SiPh chip. This transmitter is fully compatible with standard 300-mm CMOS fabrication processes. The microdisk EOIM has an electro-optic bandwidth from 19 GHz to 28 GHz, an extinction ratio (ER) of 9.8-dB, a modulator insertion loss (IL) of 2.6-dB, and an optical modulation amplitude (OMA) of 0.49 at a 2-V voltage swing. By implementing the transmitter in a high-speed optical transmission link, we demonstrates an on-off keying (OOK) data transmission at 20 Gbt/s, yielding a total transmitter capacity of 1.28 Tbt/s with a low dynamic modulation energy consumption of 29 fJ/bit. Furthermore, we utilize a home-built laser annealing system to implement automated in-situ, precise, and non-reversible post-fabrication trimming of the microdisk EOIM transmitter, and achieves a controlled minimum wavelength red shift of 100 pm. By employing this non-volatile laser trimming technique, we experimentally demonstrate a thermal-tuning free, fully passive 5-channel DWDM (50 GHz channel spacing) silicon microdisk link. Compared with the untrimmed microdisk EOIM, The developed trimming technique reduces 66 % of the ring redundancy required to achieve the fully passive WDM transmission and needs 67 % percent of the thermal energy consumption required to tune the resonant wavelength. Our non-volatile trimming approach highlights the potential of SiPh in leveraging light as a high-bandwidth and energy-efficient inter-chip communication medium, paving the way for addressing the critical scaling challenges for data-movement within modern data centers.
@@
Architecture of the High Parallelism Microdisk Transmitter
The microdisk transmitter is specifically engineered to achieve low energy consumption while simultaneously delivering high per-fiber bandwidth through WDM and substantial spatial parallelism through SDM. Two critical facts underlie the transmitter’s low energy consumption. First, the resonant microdisk cavities featuring strong optical mode confinement within the p–n depletion region. This architecture ensures a high electro-optic modulation efficiency, lowering the dynamic energy required to encode digital signals. Second, we introduce a laser trimming technology capable of permanently tuning the resonant wavelengths of microdisk modulators with picometer-scale precision. By reducing the continuous power expenditure associated with conventional thermal tuning, this approach reduces the transmitter’s static energy consumption.
Beyond energy efficiency, the transmitter incorporates a highly parallelized optical architecture that exploits both WDM and SDM simultaneously to scale up the transmission capacity. Figure 1 illustrates the schematic overview of our transmitter chip. An inverse-taper edge coupler at the chip input efficiently couples external laser illumination from a standard single-mode fiber. Subsequently, a Y-junction-based optical splitting tree uniformly distributes input optical power among 64 distinct spatial channels. Integrated thermal phase shifters, positioned along these channels, provide precise and independent control of optical phases, crucial for coherent and phase-sensitive optical modulation in densely multiplexed systems. At the heart of the transmitter, a compact array of microdisk modulators, each with a radius of only 4 µm, encodes electrical data onto optical carriers via resonance-based intensity modulation. Each microdisk EOIM features dual optical outputs strategically designed to maximize parallelism and versatility. One output connects directly to an individual grating coupler, providing isolated access to single-channel operation and thereby maximizing spatial parallelism. The second output path integrates into a combined grating coupler array, enabling high spectral density multiplexed WDM transmission. This dual-output configuration provides simultaneous flexibility in accessing discrete optical channels individually and transmitting multiplexed signals with high bandwidth density.
Critically, our parallelized modulator layout overcomes some limitations of the cascaded resonant architectures, such as progressive optical power loss and cumulative thermal and optical crosstalk. By distributing power evenly and arranging modulators in independent optical paths, our design ensures uniform modulation performance across all channels and enables precise, independent tuning and operation.
In summary, our highly parallelized silicon microdisk transmitter architecture, combining permanent wavelength trimming, and dual-output multiplexing, offers improvements in power efficiency, scalability, operational uniformity, and system flexibility. These advantages position our transmitter technology as a compelling candidate for next-generation high-capacity optical interconnects and photonic computing applications.
@@
To evaluate and analyze the performance of the transmitter, we first performed the measurement on a single-channel microdisk EOIM. Fig. 2 shows the measured transmission spectrum of the microdisk EOIM. The average off-resonance IL is 0.6 dB, with a free spectrum range (FSR) of 25.6 nm, a full width at half maximum (FWHM) of 0.29 nm/37 GHz at 0-V d.c. voltage. The apodized TE mode grating coupler has a 3-dB bandwidth of 20 nm. The modulator work in the carrier depletion mode with a reverse biased voltage. Compared with normal lateral junctions, the vertical p–n junction microdisk structure enables a low driven voltage by featuring a higher overlap between the electrical p–n depletion region and the microdisk optical whispering gallery mode. The corresponding resonant wavelengths versus the reverse biased voltage are plotted in the Fig. 2(d), showing a d.c. response among all these 64 channels with an ER of 9.8-dB, a modulator IL of 2.6-dB, and an OMA of 0.49 over a 2-V swing voltage.
The microdisk operates in the under-coupled regime, where the external coupling loss is smaller than the round-trip intrinsic loss. Increasing the reverse-biased voltage further depletes the free carriers, thereby reducing free-carrier absorption and lowering the intrinsic loss. This shift pushes the device towards critical coupling, deepening the on-resonance transmission dip and slightly increasing the ER and OMA. Owing to the strong overlap between the optical mode and the p–n depletion region, the microdisk exhibits a large electro-optic tuning efficiency of $90\leavevmode\nobreak\ \mathrm{pm\,V^{-1}}$. The figure of merit for modulation efficiency, $V_{\pi}\!\cdot\!L$, is therefore $V_{\pi}\!\cdot\!L=3.57\,\mathrm{V\!\cdot\!mm}$. By fitting the ring-resonator transmission Lorentz model, the propagation loss is found to decrease from $\alpha=8.73\leavevmode\nobreak\ \mathrm{dB\,mm^{-1}}$ to $\alpha=6.63\leavevmode\nobreak\ \mathrm{dB\,mm^{-1}}$ as the reverse bias is set from $0.5\leavevmode\nobreak\ \mathrm{V}$ to $1.5\leavevmode\nobreak\ \mathrm{V}$ (Fig. 2(f)), and the quality factor $Q$ rises from 5290 to 6510. We didn’t further increase the reverse biased voltage to avoid any potential damages. The transmission spectrum of the integrated thermal phase shifter, measured with heater voltages from $0\leavevmode\nobreak\ \mathrm{V}$ to $5\leavevmode\nobreak\ \mathrm{V}$, is plotted in Fig. 2(e). As the drive power increases, the resonance red-shifts. The phase shifter resistance is $2.0\leavevmode\nobreak\ \mathrm{k}\Omega$, the calculated thermo-optic efficiency and thermal $\pi$-phase shift powers are $420\leavevmode\nobreak\ \mathrm{pm\,mW^{-1}}$ and $30.5\leavevmode\nobreak\ \mathrm{mW}$, respectively.
@@
The EO bandwidth of the microdisk EOIM is determined by the RC time constant of the junction, the photon lifetime of the microdisk resonator, as well as the offset of the laser wavelength and the resonant wavelength. The RC determined bandwidth is extracted from the measured $S_{11}$ results. A 40.5 GHz vector network analyzer was used for the S-parameter measurements. The equivalent circuit parameters can be derived through the $S_{11}$ fitting. In the circuit, $C_{pn}$ denotes the p–n junction capacitance, $R_{pn}$ is the p–n junction series resistance, $C_{box}$ is the buried-oxide capacitance, $R_{sub}$ stands for the silicon substrate resistance, and $C_{pad}$ is the capacitance of the pads through the air. The vertical p–n junction has a fitted capacitance $C_{total}$ of 112.5 fF. Utilizing the equivalent circuits with a $50\,\Omega$ source impedance, the frequency response across the p–n diode capacitance can be plotted, which corresponds to the small-signal response that is actually effective for the modulation, as shown in Fig. 3(b), the RC-time-limited bandwidth of the modulator is $f_{\mathrm{rc}}$= 42 GHz. The second EO-bandwidth contributor, the photon lifetime of the microdisk, is characterized by the $Q$ factor of the resonant cavity. The photon-lifetime-limited bandwidth, $f_{\mathrm{ph}}$, then can be expressed by $f_{\mathrm{ph}}=\sqrt{\sqrt{2}-1}\,\dfrac{c\,\Delta\lambda}{\lambda^{2}}$ [27], where $c$ is the light speed, and $\lambda$ is the resonance wavelength. The $Q$ factor of the microdisk is 5290, so that the $f_{\mathrm{ph}}$ is around 23 GHz. The final E/O bandwidth of the microring can be calculated using $f_{\mathrm{E/O}}=\dfrac{f_{\mathrm{rc}}\,f_{\mathrm{ph}}}{\sqrt{f_{\mathrm{rc}}^{2}+f_{\mathrm{ph}}^{2}}}$. The fitted $S_{21}$ is represented by the red and green lines.
@@
The measured electro-optic response of the modulator is shown in Fig. 3(b). The laser was aligned off resonance at $\lambda=1{,}536.45\leavevmode\nobreak\ \text{nm}$ to preserve linearity and to achieve a faster response by eliminating the photon lifetime influential. A $50\leavevmode\nobreak\ \text{mV V}_{\mathrm{pp}}$ sinusoidal small signal was applied to the microdisk modulator at frequencies spanning from $50\leavevmode\nobreak\ \text{MHz}$ to $40.5\leavevmode\nobreak\ \text{GHz}$. The through-port power was measured with an external high-speed photodetector. The electro-optic 3 dB bandwidth was experimentally measured (orange and blue dash lines) from $19\leavevmode\nobreak\ \text{GHz}$ to $28\leavevmode\nobreak\ \text{GHz}$ among different channels.
@@
Figure 4: Non-volatile trimming and DWDM of the Microdisk Transmitter. (a) Experimental setup of the non-volatile, close-looped, automatic laser trimming. (b) Microdisk resonance wavelength shift versus the trimming time. (e) Wavelength distribution of the 64 channel microdisk transmitter, showing a wavelength span of around 2-nm. (D) Passive 50 GHz DWDM transmitter achieved through the laser trimming, only odd channels were selected to avoid the channel crosstalk. (e) The design redundancy channels needed to achieve the fully passive WDM for pre and post trimmed devices. (f-j) Measured 20 Gbt/s eye-diagram of the trimmed microdisk modulator, which demonstrated a fully passive, dense wavelength division (50 GHz) channel, scalar bar: 40 ps.
@@
High-speed optical eye diagrams, presented in Fig. 3(c), were obtained utilizing a high-speed electro–optic measurement setup with a real-time oscilloscope at data rates of 20  Gbt/s (setup-limited). The open eye diagram indicates a good signal integrity and robust modulation performance. The 64-channel photonic transmitter demonstrates an high bandwidth of 1.2  Tbt/s while maintaining a low dynamic energy consumption of approximately $29\leavevmode\nobreak\ \text{fJ/bit}$ per channel, calculated using the conventional $CV^{2}/4$ metric with a 1-V swing voltage. It is worth noting that 20 Gbt/s eyes are the upper limit of the experimental setup, a higher-bandwidth setup can further improve the transmitter capacity.
@@
To quantitatively evaluate the modulation accuracy, we performed comprehensive encoding and decoding characterization of the modulator. Data signals generated by the AWG were first pulse-amplitude-modulated to $2\leavevmode\nobreak\ \mathrm{Gbt/s}$ and then encoded onto the EOIM, transmitted through optical fiber, amplified by an EDFA, detected by a high-speed photodetector, and digitized using a high-speed oscilloscope. The modulation accuracy was quantitatively assessed by calculating the standard deviation of the difference between the theoretically generated AWG data and the measured oscilloscope waveforms. The measured standard deviation was $0.007$, corresponding to an effective bit precision of $7\leavevmode\nobreak\ \text{bits}$. This low standard deviation highlights the transmitter’s high modulation accuracy and stability, confirming its suitability for high-performance optical communication and computing systems.
Non-volatile trimming and passive DWDM in the Microdisk Transmitter

### v2

@@ Abstract
Optical interconnects are the most promising solution to address the data-movement bottleneck in data centers. Silicon microdisks, benefiting from their compact footprint, low energy consumption, and wavelength division multiplexing (WDM) capability, have emerged as an attractive and scalable platform for optical modulation. However, microdisk resonators inherently exhibit low fabrication error tolerance, limiting their practical deployment. Here, utilizing a CMOS photonics platform, we demonstrate 1.28 Tb/s of off-die bandwidth through a 64 microdisk modulator system. In addition, we develop an automated, close-looped, non-reversible, low-loss, and picometer-precision permanent wavelength tuning technique using laser trimming, and achieve a fully passive, 5-channel dense wavelength division multiplexing (DWDM, 50 GHz spacing) transmitter. The integration of the high speed (1.28 Tb/s), low energy consumption (29 fJ/bit) and the permanent wavelength trimming lays a robust foundation for next-generation optical interconnect systems, poised to facilitate scaling of future AI and computing hardware.
@@ Introduction
The exponential growth of artificial intelligence (AI) and deep learning applications drives computational workloads from individual devices to server racks in data centers, placing unprecedented requirements on the interconnect and communication infrastructures, which need to be high-speed and low-energy-consumption [1, 2, 3, 4, 5, 6, 7, 8]. Traditional intra- and inter- chip interconnects, which rely on electrical metal-wire links, are struggling to satisfy the increasing bandwidth and energy consumption requirements imposed by current data-hungry AI applications [3, 9, 2]. As AI workloads scale, multi-node, multi-processor systems demand significantly much higher-speed and more energy-efficient XPU-to-XPU data communications to effectively handle large-scale computational-intensive tasks [9, 10, 11, 12]. In contrast, optical interconnect technologies that leverage silicon photonics (SiPh) offer bandwidth and energy efficiency advantages over conventional electrical metal-wire links [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. SiPh components, which are based on the total internal reflection of silicon-on-insulator (SOI) waveguides, have several distinct advantages, including compatibility with mature complementary metal–oxide–semiconductor (CMOS) manufacturing process, large bandwidth (10’s THz), compact footprint ($\penalty 10000\ \upmu\text{m}$ scale), low loss (dB/cm to dB/m), and intrinsic high parallelism [27, 28, 29, 30, 31, 32, 33]. These combined advantages make SiPh interconnects an essential infrastructure components to address the large-bandwidth and low-energy-consumption requirements for contemporary data center.
@@
For short-reach interconnects and communications, the low-complexity and low-cost intensity-modulated direct-detection (IM-DD) system remains dominant within the data centers [34, 16, 17], where integrated electro-optic intensity modulators (EOIMs) convert the XPU electrical signals to analog optical intensities. The high-index contrast of the SiPh platform enables the realization of ultra-compact vertical-type p–n junction microdisk EOIMs. The vertical junction microdisk modulators provide large overlap between carriers and the optical mode, in addition, by leveraging the intrinsic wavelength selectivity nature, WDM and spatial division ultiplexing (SDM) can be implemented to provide a scalable high parallel modulation across a broadband wavelength range. Consequently, SiPh microdisk EOIMs enables large-scale, high bandwidth, high efficiency, and low energy consumption modulation in a compact footprint [29, 27, 19, 17, 35]. The CMOS compatibility of SiPh also allows high-yield mass production, which is desired for low-cost optical interconnects and communications in data centers.
@@
Figure 1: Artistic vision and microscope image of the high parallelism, high capacity micro-disk transmitter system. (a) Schematic diagram of the 64-channel high parallelism, high capacity microdisk transmitter. The transmitter supports high parallelism SDM and allows for high per-fiber bandwidth WDM transmission. (b) Optical microscope image of the transmitter chip, from input to output, the chip includes the input grating coupler array, microdisk modulators, phase shifters, y-splitter combiner tree, and output edge coupler. (c) High resolution microscope images of the main components of the chip. (d) 2D emission image of the chip shows an 8 by 8 emitter array.
@@
Figure 2: Direct current characterization of the Microdisk Transmitter chip. (a) Experimental setup of the microdisk transmitter direct current characterization, the TLS and OSA are synchronized to provide high precision characterization, an IR camera is employed to provide parallel characterization. TLS, tunable laser; OSA, optical spectrum analyzer; PC, polarization controller; VGFA, V-groove fiber array; EDFA, Er-doped fiber amplifier. (b) Measured and fitted grating coupler transmission spectrum of the transmitter. (c) Measured broadband transmission spectrum of a single-channel transmitter at 0-V D.C voltage. (d) Measured transmission spectrum versus reversed junction voltage from 0.5 V to -1.5 V, under a 2-V voltage swing at 1536.38 nm, the modulator insertion loss (IL) is 2.6 dB, the extinction ratio (ER) is 9.8 dB, and the optical modulation amplitude (OMA) is -3.2 dB relative to the input optical power, these parameter values are nearly constant among these 64-channel microdisk modulators. (e) Measured transmission spectrum versus heater voltage ranging from 0 V to 5 V, showing a wavelength red shift tunning efficiency of 420 pm/mW. (f) Extracted absorption coefficient and refractive index change of the microdisk transmitter versus junction voltages from 0.5 V to -1.5 V.
@@
While microdisk EOIMs offer compact footprints, they exhibit extreme sensitivity to fabrication errors and uncertainties [35, 36, 37, 38]. Even sub-nanometer variations in waveguide dimensions can significantly shift the spectral response of these devices. Both active tuning and post-fabrication trimming methods are employed to correct the resonance uncertainties [39, 38, 40, 41]. Active tuning, commonly achieved through the thermo-optic effect, though effective, it needs continuously large power consumption, posing a serious challenge to the scalability of large-scale implementation. In contrast, post-fabrication trimming permanently modify the effective index of the optical mode and doesn’t need continuous power consumption. Established techniques include selective post-fabrication etching, deposition of phase change materials (PCMs), anodic oxidation, laser and thermal annealing [38, 42, 43, 44]. However, in standard foundry‑fabricated compact microdisk modulators the waveguides are buried beneath thick dielectric top claddings and criss‑crossed by embedded metal interconnects, thus less compatible with PCM integration, in addition, incorporate active p–n junctions cannot tolerate the elevated post‑process temperatures required by some trimming schemes. Instead, adopt a method to locally densify the oxide cladding of each modulator via laser annealing as a practical post-fabrication trimming technique [45], we observe that this process increases the mode effective refractive index of the microdisk EOIM, yielding a red shift in the resonant wavelength.
@@
In this work, we demonstrate and perform non-volatile trimming of a highly parallel, high capacity EOIM transmitter featuring 64 compact microdisks integrated onto a single SiPh chip. The microdisk EOIM has an electro-optic bandwidth from 19 GHz to 28 GHz, an extinction ratio (ER) of 9.8-dB, an insertion loss (IL) of 2.6-dB, and an optical modulation amplitude (OMA) of -3.2 dB relative to the input optical power at a 2-V voltage swing. By implementing the transmitter in a high-speed optical transmission link, we demonstrates an on-off keying (OOK) data transmission at 20 Gbt/s (setup limited), yielding a total transmitter capacity of 1.28 Tbt/s and a low dynamic modulation energy consumption of 29 fJ/bit. Furthermore, we utilize a home-built laser annealing system to implement automated in-situ, precise, and non-reversible post-fabrication trimming of the transmitter, and achieves a wavelength red shift from 100 pm to 380 pm. By employing this non-volatile laser trimming technique, we experimentally demonstrate a thermal-tuning free, fully passive 5-channel DWDM (50 GHz channel spacing) silicon microdisk link. Compared with the untrimmed microdisk EOIM, the developed trimming technique reduces 33 % consumption of the energy required to tune the resonant wavelength. Our non-volatile trimming approach highlights the potential of SiPh in leveraging light as a high-bandwidth and energy-efficient inter-chip communication medium, paving the way to addressing the critical scaling challenges for data-movement within modern data centers.
Architecture of the High Parallelism Microdisk Transmitter
Figure 1 illustrates the structure of the transmitter chip. 64 grating couplers, each addressed by a distinct wavelength, efficiently couple external lights into the chip. Subsequently, a compact array of 64 microdisk modulators, each with a radius of only 4 µm, encodes the electrical data onto optical carriers via intensity modulation. A low-loss Y-junction–based optical combining tree then merges the 64 modulated carriers into a single bus waveguide. Finally, an inverse-taper edge coupler at the chip output efficiently transfers the spectrally multiplexed signal into a standard single-mode fiber. By admitting 64 spatial inputs while assigning a unique wavelength to each, the architecture simultaneously exploits SDM at the input and WDM at the output, yielding a highly parallelized transmitter with high bandwidth density.
@@
The parallelized modulator layout overcomes some limitations of the cascaded resonant architectures, such as progressive optical power loss and cumulative thermal and optical crosstalk. By distributing power evenly and arranging modulators in independent optical paths, our design ensures uniform modulation performance across all channels and enables precise, independent tuning and operation.
@@
To evaluate and analyze the performance of the transmitter, we first performed the measurement on a single-channel microdisk EOIM. Fig. 2 shows the measured transmission spectrum of the microdisk EOIM. The average off-resonance IL is 0.6 dB, with a free spectrum range (FSR) of 25.6 nm, a full width at half maximum (FWHM) of 0.29 nm/37 GHz at 0-V d.c. voltage. The apodized TE mode grating coupler has a 3-dB bandwidth of 20 nm. The modulator work in the carrier depletion mode with a reverse biased voltage. Compared with normal lateral junctions, the vertical p-n junction microdisk structure enables a low driven voltage by featuring a higher overlap between the electrical p–n depletion region and the microdisk optical whispering gallery mode. The corresponding resonant wavelengths versus the reverse biased voltage are plotted in the Fig. 2(d), showing a d.c. response among all these 64 channels with an ER of 9.8-dB, a modulator IL of 2.6-dB, and an OMA of -3.2 dB relative to the input optical power over a 2-V swing voltage.
The microdisk operates in the under-coupled regime, where the external coupling loss is smaller than the round-trip intrinsic loss. Increasing the reverse-biased voltage further depletes the free carriers, thereby reducing free-carrier absorption and lowering the intrinsic loss. This shift pushes the device towards critical coupling, deepening the on-resonance transmission dip and slightly increasing the ER and OMA. Owing to the strong overlap between the optical mode and the p–n depletion region, the microdisk exhibits a large electro-optic tuning efficiency of $90\penalty 10000\ \mathrm{pm\,V^{-1}}$. The figure of merit for modulation efficiency, $V_{\pi}\!\cdot\!L$, is therefore $V_{\pi}\!\cdot\!L=3.57\,\mathrm{V\!\cdot\!mm}$, which is the best results ever obtained for the carrier depletion mode silicon modulators. By fitting the ring-resonator transmission Lorentz model, the propagation loss is found to decrease from $\alpha=8.73\penalty 10000\ \mathrm{dB\,mm^{-1}}$ to $\alpha=6.63\penalty 10000\ \mathrm{dB\,mm^{-1}}$ as the reverse bias is set from $0.5\penalty 10000\ \mathrm{V}$ to $1.5\penalty 10000\ \mathrm{V}$ (Fig. 2(f)), and the quality factor $Q$ rises from 5290 to 6510. We didn’t further increase the reverse biased voltage to avoid any potential damages. The transmission spectrum of the integrated thermal phase shifter, measured with heater voltages from $0\penalty 10000\ \mathrm{V}$ to $5\penalty 10000\ \mathrm{V}$, is plotted in Fig. 2(e). As the drive power increases, the resonance red-shifts. The phase shifter resistance is $2.0\penalty 10000\ \mathrm{k}\Omega$, the calculated thermo-optic efficiency and thermal $\pi$-phase shift powers are $420\penalty 10000\ \mathrm{pm\,mW^{-1}}$ and $30.5\penalty 10000\ \mathrm{mW}$, respectively.
@@
The EO bandwidth of the microdisk EOIM is determined by the RC time constant of the junction, the photon lifetime of the microdisk resonator, as well as the offset of the laser wavelength and the resonant wavelength. The RC determined bandwidth is extracted from the measured $S_{11}$ results. A 40.5 GHz vector network analyzer was used for the S-parameter measurements. The equivalent circuit parameters can be derived through the $S_{11}$ fitting. In the circuit, $C_{pn}$ denotes the p–n junction capacitance, $R_{pn}$ is the p–n junction series resistance, $C_{box}$ is the buried-oxide capacitance, $R_{sub}$ stands for the silicon substrate resistance, and $C_{pad}$ is the capacitance of the pads through the air. The vertical p–n junction has a fitted capacitance $C_{total}$ of 112.5 fF. Utilizing the equivalent circuits with a $50\,\Omega$ source impedance, the frequency response across the p–n diode capacitance can be plotted, which corresponds to the small-signal response that is actually effective for the modulation, as shown in Fig. 3(b), the RC-time-limited bandwidth of the modulator is $f_{\mathrm{rc}}$= 42 GHz. The second EO-bandwidth contributor, the photon lifetime of the microdisk, is characterized by the $Q$ factor of the resonant cavity. The photon-lifetime-limited bandwidth, $f_{\mathrm{ph}}$, then can be expressed by $f_{\mathrm{ph}}=\sqrt{\sqrt{2}-1}\,\dfrac{c\,\Delta\lambda}{\lambda^{2}}$ [27], where $c$ is the light speed, and $\lambda$ is the resonance wavelength. The $Q$ factor of the microdisk is 5290, so that the $f_{\mathrm{ph}}$ is around 23 GHz. The final E/O bandwidth of the microdisk can be calculated using $f_{\mathrm{E/O}}=\dfrac{f_{\mathrm{rc}}\,f_{\mathrm{ph}}}{\sqrt{f_{\mathrm{rc}}^{2}+f_{\mathrm{ph}}^{2}}}$. The fitted $S_{21}$ is represented by the red and green lines.
@@
The measured electro-optic response of the modulator is shown in Fig. 3(b). The laser was aligned off resonance at $\lambda=1{,}536.45\penalty 10000\ \text{nm}$ to preserve linearity and to achieve a faster response by eliminating the photon lifetime influential. A $50\penalty 10000\ \text{mV V}_{\mathrm{pp}}$ sinusoidal small signal was applied to the microdisk modulator at frequencies spanning from $50\penalty 10000\ \text{MHz}$ to $40.5\penalty 10000\ \text{GHz}$. The through-port power was measured with an external high-speed photodetector. The electro-optic 3 dB bandwidth was experimentally measured (orange and blue dash lines) from $19\penalty 10000\ \text{GHz}$ to $28\penalty 10000\ \text{GHz}$ among different channels.
@@
Figure 4: Non-volatile trimming of the Microdisk Transmitter. (a) Experimental setup of the non-volatile, close-looped, automatic laser trimming, the 2D grating array was designed to provide in-time monitoring of the emission wavelength. (b) Microdisk resonance wavelength shift versus the trimming time, showing a minimum wavelegnth red shift of 100 pm. (c) Thermal tunning energy saving versus the trimming range. (d) The design redundancy channels needed to achieve the fully passive WDM for pre and post trimmed devices.
@@
High-speed optical eye diagrams, presented in Fig. 3(c), were obtained utilizing a high-speed electro–optic measurement setup with a real-time oscilloscope at data rates of 20 Gbt/s (setup-limited). The open eye diagram indicates a good signal integrity and robust modulation performance. The 64-channel photonic transmitter demonstrates an high bandwidth of 1.28 Tbt/s while maintaining a low dynamic energy consumption of approximately $29\penalty 10000\ \text{fJ/bit}$ per channel, calculated using the conventional $CV^{2}/4$ metric with a 1-V swing voltage. It is worth noting that 20 Gbt/s eyes are the upper limit of the experimental setup, a higher-bandwidth setup can further improve the transmitter capacity [17].
@@
To quantitatively evaluate the modulation accuracy, we performed comprehensive encoding and decoding characterization of the modulator. Data signals generated by the AWG were first pulse-amplitude-modulated to $2\penalty 10000\ \mathrm{Gbt/s}$ and then encoded onto the EOIM, transmitted through optical fiber, amplified by an EDFA, detected by a high-speed photodetector, and digitized using a high-speed oscilloscope. The modulation accuracy was quantitatively assessed by calculating the standard deviation of the difference between the theoretically generated AWG data and the measured oscilloscope waveforms. The measured standard deviation of the difference was $0.007$, corresponding to an effective bit precision of $7\penalty 10000\ \text{bits}$. This low standard deviation highlights the transmitter’s high modulation accuracy and stability, confirming its suitability for high-performance optical communication and computing systems.
Non-volatile trimming of the Microdisk Transmitter
Figure 5: Fully passive DWDM transmission. (a) Wavelength distribution of the foundry fabricated microdisk chip. (b) Selected DWDM wavelength distribution after the trimming, showing uniform wavelength spacing of 25 GHz. (c-h) Schematic and results of the 20 Gbt/s eye-diagram of the trimmed microdisk modulators, which demonstrated a fully passive, dense wavelength division (50 GHz) channel, scalar bar: 40 ps.